(92) Statistics. Sam must have an average of 70 or more in his summer course to obtain a grade of C. His first three test grades were 75,63, and 68. Write an inequality representing the score that Sam must get on the last test to get a C grade.
(94) Business and finance. The cost for a long-distance telephone call is $0.36 for the first minute and $0.21 for each additional minute or portion thereof. Write an inequality representing the number of minutes a person could talk without exceeding $3.



Answer by venugopalramana(3286)

(92) Statistics. Sam must have an average of 70 or more in his summer course to obtain a grade of C. His first three test grades were 75,63, and 68. Write an inequality representing the score that Sam must get on the last test to get a C grade.
TOTAL MARKS IN 3 TESTS = 75+63+68=206
LET SCORE IN THE LAST TEST = X
TOTAL SCORE IN 4 TESTS = 206+X
AVERAGE IN 4 TESTS = A = (206+X)/4
A>=70 HENCE
(206+X)/4 >=70
206+X >= 4*70=280
X>= 280-206=74
(94) Business and finance. The cost for a long-distance telephone call is $0.36 for the first minute and $0.21 for each additional minute or portion thereof. Write an inequality representing the number of minutes a person could talk without exceeding $3.
L.D.CALL CHARGE FOR 1 ST.MINUTE = 0.36
LET THE TOTAL NUMBER OF MINUTES TALKED = T MTS.
ADDITIONAL MINUTES TALKED OVER THE I MINUTE = T-1
CHARGES FORT ADDITIONAL MINUTES = (T-1)*0.21
TOTAL CHARGE = 0.36+(T-1)*0.21<=3.00 $
0.36+0.21T-0.21<=3
0.21T<=3-0.36+0.21=2.85
T<=2.85/0.21=285/21=95/7.